# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v3

@@ 5.1.1 Experimental Setup
Tasks and datasets. We evaluate our algorithm under three types of dynamics shifts, namely morphology, kinematic and friction changes, across three OpenAI Gym locomotion tasks: HalfCheetah, Hopper, and Walker2d [brockman2016openai]. Each experiment involves an offline environment and an online environment with modified transition dynamics following [lyu2024odrl]. Morphology shifts alter the sizes of body parts, kinematic shifts impose constraints on joint angles, and friction shifts modify the static, dynamic, and rolling friction coefficients. For each task, we use three D4RL source datasets: medium, medium replay, and medium expert, which capture varying levels of data quality [fu2020d4rl]. Additional environment details are in Appendix K.
Baselines. We compare CompFlow against the following baselines: BC-SAC extends SAC by incorporating both offline and online data, with a behavior cloning (BC) term for the offline data. H2O [niu2022trust] penalizes Q-values for state-action pairs with large dynamics gaps. BC-VGDF [xu2023cross] selects offline transitions with value targets consistent with the online environment and adds a BC term. BC-PAR [lyu2024cross] applies a reward penalty based on representation mismatch between offline and online transitions, also including a BC term. Implementation details are provided in Appendix K.
@@ 5.1.2 Main Results
We evaluate the return of each algorithm in the target domain after 40K environment interactions and 400K gradient steps, reflecting the limited interaction setting. The performance of all methods is reported in Table 1. Our key findings are summarized as follows:
(1) CompFlow consistently outperforms recent off-dynamics reinforcement learning baselines across a wide range of dataset qualities and types of dynamics shifts. Specifically, it achieves the highest return on 19 out of 27 tasks and matches the top performance on an additional 5 tasks. On average, CompFlow attains a score of 2193, compared to 1920 achieved by the second-best baseline, BC-SAC, representing a relative improvement of 14.2%. To further validate the robustness of our method, we present additional experimental results under even more pronounced dynamics shifts in Appendix J, which demonstrates consistent findings.
(2) Incorporating offline data can significantly enhance learning performance. On average, CompFlow achieves a 149.8% improvement over SAC. In addition, CompFlow outperforms the base algorithm BC-SAC, which directly incorporates all offline data, on 23 out of 27 tasks and matches its performance on the remaining 3, providing strong empirical support for the effectiveness of our method.
(3) Additionally, we observe that recent baselines often perform similarly to BC-SAC across many tasks, consistent with the findings in [lyu2024odrl], suggesting that they fail to effectively leverage the offline source domain data. This may stem from methods like H2O and BC-PAR relying on KL divergence or mutual information, which can be ill-defined under large dynamics gaps or different supports. Meanwhile, BC-VGDF adopts a data filtering strategy based on estimated value functions, which is often more challenging than learning transition dynamics due to bootstrapping bias and target non-stationarity. In contrast, our method, CompFlow, leverages flow matching to model complex transition dynamics and exploits its theoretical connection to optimal transport to estimate the dynamics gap via Wasserstein distance—a more robust and principled metric.
@@ 5.1.3 Ablation and Hyperparameter Analysis
Composite Flow. We first evaluate the effectiveness of the proposed composite flow matching design. For comparison, we train a direct flow model on the target domain initialized from a Gaussian prior. We compute the mean squared error (MSE) on a held-out 10% validation set and report the average MSE across different epochs during RL training. As shown in Figure 3, the composite flow significantly reduces the MSE of the transition dynamics on the target domain. This improvement stems from its ability to reuse structural knowledge learned from the source domain, where abundant data is available. These empirical findings support the theoretical insight presented in Theorem 3.1 that our composite flow can improve the generalization ability.
Impact of data selection ratio $\xi\%$. The data selection ratio $\xi$ decides how many source domain data in a sampled batch can be shared for policy training. A larger $\xi$ indicates that more source domain data will be admitted. To examine its influence, We sweep $\xi$ across {70, 50, 30, 20}. The results is shown in Figure 4. Although the optimal $\xi$ seems task dependent, moderate values of $\xi$ (e.g., 30 or 50) generally yield good performance across tasks, striking a balance between leveraging useful source data and avoiding high dynamics mismatch. When $\xi$ is too large (e.g., 70), performance often degrades, particularly in Walker Medium (Morph) and Walker Medium Replay (Morph), likely due to the inclusion of low-quality transitions with large dynamics gap. HalfCheetah consistently achieves higher returns compared to other domains, suggesting that knowledge transfer and policy learning in this environment are easier. Consequently, overly conservative filtering (e.g., $\xi=20$) may exclude valuable source data, leading to slower learning. In this case, allowing more source data (e.g., $\xi=70$) appears beneficial.
@@ 5.2 Patrol Policy Learning for Wildlife Conservation
We evaluate CompFlow in a wildlife conservation setting, where the objective is to learn a patrol policy that allocates limited ranger effort to reduce poaching and protect wildlife. We adopt the simulation environment introduced by xu2021robust. The environment is represented as a spatial grid, partitioned into $1\times 1$ km cells. At each time step, the agent selects a distribution of patrol effort across the cells, subject to a fixed budget. The state, which includes wildlife population density and poaching risk for each cell, evolves over time according to dynamics that depend on both poacher behavior and the deployed patrols. The per-step reward reflects the expected number of animals preserved through deterrence and spatial coverage achieved by the patrol strategy.
@@
We assume access to an offline dataset collected under a previous policy in Murchison Falls National Park. The objective is to use this data to learn an improved patrol policy for Queen Elizabeth National Park, with only limited interaction with the new environment. Due to differences in ecological conditions and poacher behavior, the transition dynamics in the two parks are not the same. Figure 6 presents the results. CompFlow achieves the highest reward, exceeding the best baseline, BC-PAR, by 8.8%. Compared to training a policy from scratch using SAC, CompFlow improves the average reward by 20.8% under the same interaction budget. This improvement is especially important in conservation settings where ranger capacity is limited and the protected area is large. For instance, Murchison Falls National Park spans approximately 3,900 square kilometers, while Queen Elizabeth National Park covers about 1,980 square kilometers. The large size of these areas makes full coverage difficult, and learning effective patrol strategies with minimal exploration is crucial for protecting wildlife.
@@ Appendix A Limitations
(1) CompFlow is currently limited to settings where the offline data and the online environment share the same state and action spaces. Future work could explore estimating the dynamics gap in a shared latent embedding space to relax this assumption [zhuang2023dygen]. (2) Our evaluation is conducted entirely in simulated environments; applying CompFlow to real-world scenarios remains an important direction for future work. (3) We have not yet incorporated the uncertainty in estimating the dynamics gap. Future work may explore how to quantify this uncertainty and use it to improve both data filtering and exploration strategies [kong2023uncertainty, li2024muben].
@@ Appendix C Algorithms of Training Optimal Transport Flow Matching
Algorithm 1 Training Optimal Transport Flow Matching
1:Dataset $\mathcal{D}$, batch size $k$, cost function $c(\cdot,\cdot)$, learning rate $\mathsf{lr}$, flow model $v_{\theta}$
2:while not converged do
3:  Sample latent batch $\{x_{0}^{(i)}\}_{i=1}^{k}\sim\mathcal{N}(0,\mathbf{I})$
4:  Sample data batch $\{x_{1}^{(j)}\}_{j=1}^{k}\sim\mathcal{D}$
5:  Compute optimal transport plan:
[…]
$\triangleright$ $B_{k}$: set of $k\times k$ doubly-stochastic matrices
6:  Sample $k$ index pairs $\{(i_{\ell},j_{\ell})\}_{\ell=1}^{k}\sim A(i,j)$
7:  Sample interpolation times $\{t^{(\ell)}\}_{\ell=1}^{k}\sim\mathcal{U}[0,1]$
8:  for $\ell=1$ to $k$ do
9:   Define matched pair:
$x_{0}^{(\ell)}\doteq x_{0}^{(i_{\ell})},\quad x_{1}^{(\ell)}\doteq x_{1}^{(j_{\ell})}$
10:   Interpolate:
$x_{t}^{(\ell)}\doteq t^{(\ell)}x_{1}^{(\ell)}+(1-t^{(\ell)})x_{0}^{(\ell)}$
11:  end for
12:  Compute loss:
$\mathcal{L}\doteq\frac{1}{k}\sum_{\ell=1}^{k}\left\|v_{\theta}(x_{t}^{(\ell)},t^{(\ell)})-(x_{1}^{(\ell)}-x_{0}^{(\ell)})\right\|_{2}^{2}$
13:  Update parameters:
$\theta\leftarrow\theta-\mathsf{lr}\nabla_{\theta}\mathcal{L}$
14:end while
15:return $\theta$
@@ Appendix D Algorithms of Training Offline and Online Flows
Algorithm 2 Training of Offline Flow Matching
1:Offline dataset $\mathcal{D}_{\text{off}}$, batch size $k$, flow model $v_{\theta}$, learning rate $\mathsf{lr}$
2:while not converged do
3:  Sample a batch of transitions $\{(s^{(i)},a^{(i)},s^{\prime(i)})\}_{i=1}^{k}\sim\mathcal{D}_{\text{off}}$
4:  For each $i$, define target sample $x_{1}^{(i)}\doteq s^{\prime(i)}$
5:  Sample $\{x_{0}^{(i)}\}_{i=1}^{k}\sim\mathcal{N}(0,\mathbf{I})$
6:  Sample interpolation times $\{t^{(i)}\}_{i=1}^{k}\sim\mathcal{U}[0,1]$
7:  for $i=1$ to $k$ do
8:   Compute interpolated point: $x_{t}^{(i)}=t^{(i)}x_{1}^{(i)}+(1-t^{(i)})x_{0}^{(i)}$
9:  end for
10:  Compute training loss:
$\mathcal{L}=\frac{1}{k}\sum_{i=1}^{k}\left\|v_{\theta}(x_{t}^{(i)},t^{(i)},s^{(i)},a^{(i)})-(x_{1}^{(i)}-x_{0}^{(i)})\right\|_{2}^{2}$
11:  Update $\theta\leftarrow\theta-\mathsf{lr}\nabla_{\theta}\mathcal{L}$
12:end while
13:return Trained parameters $\theta$
Algorithm 3 Training of Online Flow via Optimal Transport
1:Pre-trained offline flow model $\psi_{\theta}(x,t\mid s,a)$, offline dataset $\mathcal{D}_{\text{off}}$, target dataset $\mathcal{D}_{\text{on}}$, batch size $k$, cost function $c(\cdot)$, regularization weight $\eta$, learning rate $\mathsf{lr}$, target flow model $v_{\phi}$
2:while not converged do
3:  Sample latent vectors $\{x_{0}^{(i)}\}_{i=1}^{k}\sim\mathcal{N}(0,\mathbf{I})$
4:  Sample offline batch $\{s_{\rm off}^{(i)},a_{\rm off}^{(i)}\}_{i=1}^{k}\sim\mathcal{D}_{\rm off}$
5:  for $i=1$ to $k$ do
6:   Predict next state for the offline environment:
$s^{\prime(i)}_{\rm off}\doteq x_{1}^{(i)}\doteq\psi_{\theta}(x_{0}^{(i)},1\mid s_{\rm off}^{(i)},a_{\rm off}^{(i)})$
7:  end for
8:  Sample online batch $\{s_{\rm on}^{(j)},a_{\rm on}^{(j)},s^{\prime(j)}_{\rm on}\}_{j=1}^{k}\sim\mathcal{D}_{\rm on}$
9:  Define $x_{2}^{(j)}\doteq s^{\prime(j)}_{\rm on}$
10:  Compute optimal transport plan:
$A\doteq\arg\min_{A\in B_{k}}\sum_{i,j}A(i,j)\cdot c(i,j)$
11:  where the cost is:
$c(i,j)\doteq\|x_{1}^{(i)}-x_{2}^{(j)}\|_{2}^{2}+\eta\left(\|s_{\rm off}^{(i)}-s_{\rm on}^{(j)}\|^{2}_{2}+\|a_{\rm off}^{(i)}-a_{\rm on}^{(j)}\|^{2}_{2}\right)$
$\triangleright$ $B_{k}$: set of $k\times k$ doubly-stochastic matrices
12:  Sample $k$ index pairs $\{(i_{\ell},j_{\ell})\}_{\ell=1}^{k}\sim A(i,j)$
13:  Sample interpolation times $\{t^{(\ell)}\}_{\ell=1}^{k}\sim\mathcal{U}[1,2]$ $\triangleright$ Time $t$ for the online flow is from 1 to 2
14:  for $\ell=1$ to $k$ do
15:   Compute:
$x_{t}^{(\ell)}\doteq(t^{(\ell)}-1)x_{2}^{(i_{\ell})}+(2-t^{(\ell)})x_{1}^{(j_{\ell})}$
16:  end for
17:  Compute training loss:
$\mathcal{L}\doteq\frac{1}{k}\sum_{\ell=1}^{k}\left\|v_{\phi}(x_{t}^{(\ell)},t^{(\ell)},s_{\rm on}^{(\ell)},a_{\rm on}^{(\ell)})-(x_{2}^{(j_{\ell})}-x_{1}^{(i_{\ell})})\right\|_{2}^{2}$
18:  Update parameters:
$\phi\leftarrow\phi-\mathsf{lr}\nabla_{\phi}\mathcal{L}$
19:end while
20:return $\phi$

### v4

@@ 5.1.1 Experimental Setup
Tasks and datasets. We evaluate our algorithm under three types of dynamics shifts, namely morphology, kinematic and friction changes, across three OpenAI Gym locomotion tasks: HalfCheetah, Hopper, and Walker2d [4]. Each experiment involves an offline environment and an online environment with modified transition dynamics following [38]. Morphology shifts alter the sizes of body parts, kinematic shifts impose constraints on joint angles, and friction shifts modify the static, dynamic, and rolling friction coefficients. For each task, we use three D4RL source datasets: medium, medium replay, and medium expert, which capture varying levels of data quality [16]. Additional environment details are in Appendix K.
Baselines. We compare CompFlow against the following baselines: BC-SAC extends SAC by incorporating both offline and online data, with a behavior cloning (BC) term for the offline data. H2O [46] penalizes Q-values for state-action pairs with large dynamics gaps. BC-VGDF [65] selects offline transitions with value targets consistent with the online environment and adds a BC term. BC-PAR [37] applies a reward penalty based on representation mismatch between offline and online transitions, also including a BC term. Implementation details are provided in Appendix K.
@@ 5.1.2 Main Results
We evaluate each algorithm by measuring its return in the target environment after 40K environment interactions and 400K gradient updates, reflecting a limited-interaction setting. The results are reported in Table 1. Our key findings are summarized as follows:
(1) CompFlow consistently outperforms recent off-dynamics RL baselines across diverse offline dataset qualities and dynamics shifts. It achieves the best return on 19/27 tasks and ties for best on 5 more. On average, CompFlow reaches a score of 2193, compared to 1920 for the strongest baseline (BC-SAC), a 14.2% relative improvement. Additional results under more severe dynamics shifts in Appendix J show the same trend.
(2) Leveraging offline data substantially improves performance: on average, CompFlow improves over SAC by 149.8%. Moreover, CompFlow outperforms BC-SAC, which directly uses all offline data, on 23/27 tasks and matches it on the remaining 3, highlighting the benefit of our dynamics-gap-aware design.
(3) We also find that many recent baselines perform similarly to BC-SAC, consistent with [38], suggesting limited ability to effectively exploit offline data under dynamics shift. One reason is that methods such as H2O and BC-PAR estimate the dynamics gap using KL divergence or mutual information, which can be ill-defined under large shifts or mismatched support. BC-VGDF instead filters data based on value estimates, which is often harder due to bootstrapping bias and target non-stationarity. In contrast, CompFlow models transition dynamics with flow matching and estimates the dynamics gap via the Wasserstein distance through its connection to optimal transport, yielding a more principled and robust criterion.
@@ 5.1.3 Ablation and Hyperparameter Analysis
Composite Flow. We first evaluate the effectiveness of the proposed composite flow matching design. For comparison, we train a direct flow model on the target environment initialized from a Gaussian prior. We compute the mean squared error (MSE) on a held-out 10% validation set and report the average MSE across different epochs during RL training. As shown in Figure 3, the composite flow significantly reduces the MSE of the transition dynamics in the environment. This improvement stems from its ability to reuse structural knowledge learned from the abundant offline data. These empirical findings support the theoretical insight presented in Theorem 3.1 that our composite flow can improve the generalization ability.
Impact of data selection ratio $\xi\%$. The data selection ratio $\xi$ decides how many offline data in a sampled batch can be shared for policy training. A larger $\xi$ indicates that more offline data will be admitted. To examine its influence, We sweep $\xi$ across {70, 50, 30, 20}. The results is shown in Figure 4. Although the optimal $\xi$ seems task dependent, moderate values of $\xi$ (e.g., 30 or 50) generally yield good performance across tasks, striking a balance between leveraging useful offline data and avoiding high dynamics mismatch. When $\xi$ is too large (e.g., 70), performance often degrades, particularly in Walker Medium (Morph) and Walker Medium Replay (Morph), likely due to the inclusion of low-quality transitions with large dynamics gap. HalfCheetah consistently achieves higher returns compared to other domains, suggesting that knowledge transfer and policy learning in this environment are easier. Consequently, overly conservative filtering (e.g., $\xi=20$) may exclude valuable offline data, leading to slower learning. In this case, allowing more offline data (e.g., $\xi=70$) appears beneficial.
@@ 5.2 Patrol Policy Learning for Wildlife Conservation
We evaluate CompFlow in a wildlife conservation setting, where the goal is to learn a patrol policy that allocates limited ranger effort to reduce poaching. We adopt the simulation environment of Xu et al. [67], which represents the park as a $1\times 1$ km spatial grid. At each time step, the agent allocates a distribution of patrol effort across cells under a fixed budget. The state, including wildlife density and poaching risk, evolves according to dynamics driven by both poacher behavior and patrol deployment, and the per-step reward reflects the expected number of animals preserved through deterrence and spatial coverage.
@@
We assume access to an offline dataset collected under a previous policy in Murchison Falls National Park, and aim to leverage it to learn an improved patrol policy for Queen Elizabeth National Park with only limited online interactions. Since ecological conditions and poacher behavior differ across parks, the transition dynamics are mismatched. As shown in Figure 6, CompFlow achieves the highest reward, outperforming the best baseline, BC-PAR, by 8.8%. It also improves over training from scratch with SAC by 20.8% under the same interaction budget. These gains are particularly valuable in large protected areas where ranger capacity is constrained: Murchison Falls spans roughly 3,900 km2, while Queen Elizabeth covers about 1,980 km2, making data-efficient learning essential for effective protection.
@@ Appendix A Limitations
(1) CompFlow is currently limited to settings where the offline data and the online environment share the same state and action spaces. Future work could explore estimating the dynamics gap in a shared latent embedding space to relax this assumption [71]. (2) Our evaluation is conducted entirely in simulated environments; applying CompFlow to real-world scenarios remains an important direction for future work. (3) We have not yet incorporated the uncertainty in estimating the dynamics gap. Future work may explore how to quantify this uncertainty and use it to improve both data filtering and exploration strategies [28, 33].
@@ Appendix C Algorithms of Training Optimal Transport Flow Matching
Algorithm 1 Training OT Flow Matching (OT-FM)
1:Dataset $\mathcal{D}$; batch size $k$; learning rate $\mathsf{lr}$; vector field $v_{\theta}$
2:Trained parameters $\theta$
3:while not converged do
4:  (Sample) Draw latent codes $(x_{0}^{(i)})_{i=1}^{k}\sim\mathcal{N}(0,\mathbf{I})$
5:  Draw data points $(x_{1}^{(j)})_{j=1}^{k}\sim\mathcal{D}$
6:  (OT coupling) Set $C_{ij}\leftarrow\|x_{0}^{(i)}-x_{1}^{(j)}\|_{2}^{2}$ for all $(i,j)\in[k]\times[k]$
7:  Compute an optimal transport plan
$A\leftarrow\arg\min_{A\in\mathcal{B}_{k}}\ \langle A,C\rangle,\qquad\mathcal{B}_{k}:=\{A\in\mathbb{R}_{+}^{k\times k}:A\mathbf{1}=\tfrac{1}{k}\mathbf{1},\ A^{\top}\mathbf{1}=\tfrac{1}{k}\mathbf{1}\}.$
$\triangleright$ AA is a (scaled) doubly-stochastic coupling
8:  Sample index pairs $(i_{\ell},j_{\ell})_{\ell=1}^{k}$ i.i.d. from the categorical distribution on $k]\times[k]$ with probabilities $A_{ij}$
9:  Sample times $(t_{\ell})_{\ell=1}^{k}\sim\mathcal{U}[0,1]$
10:  for $\ell=1$ to $k$ do
11:   $x_{t_{\ell}}^{(\ell)}\leftarrow(1-t_{\ell})\,x_{0}^{(i_{\ell})}+t_{\ell}\,x_{1}^{(j_{\ell})}$
12:   $\Delta^{(\ell)}\leftarrow x_{1}^{(j_{\ell})}-x_{0}^{(i_{\ell})}$
13:   $\ell_{\ell}\leftarrow\big\|v_{\theta}(x_{t_{\ell}}^{(\ell)},t_{\ell})-\Delta^{(\ell)}\big\|_{2}^{2}$
14:  end for
15:  $\mathcal{L}\leftarrow\frac{1}{k}\sum_{\ell=1}^{k}\ell_{\ell}$
16:  $\theta\leftarrow\theta-\mathsf{lr}\,\nabla_{\theta}\mathcal{L}$
17:end while
18:return $\theta$
@@ Appendix D Algorithms of Training Offline and Online Flows
Algorithm 2 Training the Offline Flow via Flow Matching
1:Offline dataset $\mathcal{D}_{\rm off}$; batch size $k$; offline vector field $v_{\theta}$; learning rate $\mathsf{lr}$
2:Trained offline flow parameters $\theta$
3:while not converged do
4:  Sample transitions $(s^{(i)},a^{(i)},s^{\prime(i)})_{i=1}^{k}\sim\mathcal{D}_{\rm off}$
5:  Set $x_{1}^{(i)}\leftarrow s^{\prime(i)}$ for all $i\in[k]$
6:  Sample latent codes $(x_{0}^{(i)})_{i=1}^{k}\sim\mathcal{N}(0,\mathbf{I})$
7:  Sample times $(t_{i})_{i=1}^{k}\sim\mathcal{U}[0,1]$
8:  for $i=1$ to $k$ do
9:   $x_{t_{i}}^{(i)}\leftarrow(1-t_{i})\,x_{0}^{(i)}+t_{i}\,x_{1}^{(i)}$
10:   $\Delta^{(i)}\leftarrow x_{1}^{(i)}-x_{0}^{(i)}$
11:   $\ell_{i}\leftarrow\big\|v_{\theta}(x_{t_{i}}^{(i)},t_{i},s^{(i)},a^{(i)})-\Delta^{(i)}\big\|_{2}^{2}$
12:  end for
13:  $\mathcal{L}\leftarrow\frac{1}{k}\sum_{i=1}^{k}\ell_{i}$
14:  $\theta\leftarrow\theta-\mathsf{lr}\,\nabla_{\theta}\mathcal{L}$
15:end while
16:return $\theta$
Algorithm 3 Training the Online Flow via Optimal Transport
1:Pre-trained offline flow sampler $\psi_{\theta}(x,t\mid s,a)$; online replay buffer $\mathcal{D}_{\rm on}$; batch size $k$; regularization weight $\eta>0$; learning rate $\mathsf{lr}$; initialized online vector field $v_{\phi}$
2:Trained online flow parameters $\phi$
3:while not converged do
4:  (Offline-synthetic batch) Sample $(s_{\rm off}^{(i)},a_{\rm off}^{(i)})_{i=1}^{k}\sim\mathcal{D}_{\rm on}$
5:  Sample latent codes $(x_{0}^{(i)})_{i=1}^{k}\sim\mathcal{N}(0,\mathbf{I})$
6:  for $i=1$ to $k$ do
7:   $x_{1}^{(i)}\leftarrow s_{\rm off}^{\prime(i)}\leftarrow\psi_{\theta}(x_{0}^{(i)},1|s_{\rm off}^{(i)},a_{\rm off}^{(i)})$
8:  end for
9:  (Online batch) Sample $(s_{\rm on}^{(j)},a_{\rm on}^{(j)},s_{\rm on}^{\prime(j)})_{j=1}^{k}\sim\mathcal{D}_{\rm on}$
10:  $x_{2}^{(j)}\leftarrow s_{\rm on}^{\prime(j)}$ for all $j\in[k]$
11:  (OT coupling) For all $(i,j)\in[k]\times[k]$, set
$C_{ij}\leftarrow\|x_{1}^{(i)}-x_{2}^{(j)}\|_{2}^{2}+\eta\big(\|s_{\rm off}^{(i)}-s_{\rm on}^{(j)}\|_{2}^{2}+\|a_{\rm off}^{(i)}-a_{\rm on}^{(j)}\|_{2}^{2}\big).$
12:  Compute an optimal transport plan
$A\leftarrow\arg\min_{A\in\mathcal{B}_{k}}\ \langle A,C\rangle,\qquad\mathcal{B}_{k}:=\{A\in\mathbb{R}_{+}^{k\times k}:A\mathbf{1}=\tfrac{1}{k}\mathbf{1},\ A^{\top}\mathbf{1}=\tfrac{1}{k}\mathbf{1}\}.$
$\triangleright$ AA is a (scaled) doubly-stochastic coupling
13:  Sample index pairs $(i_{\ell},j_{\ell})_{\ell=1}^{k}$ i.i.d. from the categorical distribution on $k]\times[k]$ with probabilities $A_{ij}$
14:  Sample times $(t_{\ell})_{\ell=1}^{k}\sim\mathcal{U}[1,2]$
15:  for $\ell=1$ to $k$ do
16:   $x_{t_{\ell}}^{(\ell)}\leftarrow(2-t_{\ell})\,x_{1}^{(i_{\ell})}+(t_{\ell}-1)\,x_{2}^{(j_{\ell})}$
17:   $\Delta^{(\ell)}\leftarrow x_{2}^{(j_{\ell})}-x_{1}^{(i_{\ell})}$
18:   $\ell_{\ell}\leftarrow\big\|v_{\phi}(x_{t_{\ell}}^{(\ell)},t_{\ell},s_{\rm on}^{(j_{\ell})},a_{\rm on}^{(j_{\ell})})-\Delta^{(\ell)}\big\|_{2}^{2}$
19:  end for
20:  $\mathcal{L}\leftarrow\frac{1}{k}\sum_{\ell=1}^{k}\ell_{\ell}$
21:  $\phi\leftarrow\phi-\mathsf{lr}\,\nabla_{\phi}\mathcal{L}$
22:end while
23:return $\phi$